OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

OpenStudy (anonymous):

"Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS. " By the phrase "NO SOLUTION" do you mean "NO UNIQUE SOLUTION"? or "NO POSSIBLE SOLUTION"? (and yes there is a difference!) If you meant "NO POSSIBLE SOLUTION", then how about this one: 3x – x = x + x + 3 Solving ... step ➊: Collect like terms ➞ 2x = 2x +3 step ➋: Subtract 2x from both sides ➞ 0 = 3 This of course is a untrue conclusion, which implies the original equation has no possible solution; there is NO number that will ever make this equation true. If you meant "NO UNIQUE SOLUTION", then how about this one: 3x – x = x + x Solving ... step ➊: Collect like terms ➞ 2x = 2x step ➋: Subtract 2x from both sides ➞ 0 = 0 This is a "true" conclusion, .... but .... what happened to "x"? This implies the original equation has no UNIQUE solution; that is to say, EVERY number that you can think of will make this equation true. This is a special kind of equation called an IDENTITY. Source: http://answers.yahoo.com/question/index?qid=20101101090101AAEtJNL